# SmoothCache: A Universal Inference Acceleration Technique for Diffusion Transformers

### v1

@@ 3.3 Ablations
We ablate our investigations by showing how model performance varies across different step sizes, and show the robustness of SmoothCache for the DiT-XL-256x256 model.
@@ 5 Conclusion
We introduce SmoothCache, a training-free caching technique that works with multiple diffusion solvers and multiple modalities. By leveraging layer-wise representation error from a calibration inference pass, SmoothCache identifies redundancies in the diffusion process, allowing it to cache and reuse output feature maps, reducing the number of computationally expensive layer operations. Our evaluations show it matches or exceeds the performance of existing modality-specific caching methods.

### v2

@@ 3.3 Ablations
We ablate our investigations by showing how model performance varies across different step sizes, and show the robustness of SmoothCache for the DiT-XL-256$\times$256 model.
@@ 5 Conclusion
We introduce SmoothCache, a training-free caching technique that works across multiple diffusion solvers and modalities. Using the layer representation error of a calibration inference pass, SmoothCache identifies redundancies in the diffusion process, enabling caching and reuse of output feature maps, reducing the number of computationally expensive layer operations. Our evaluations show it at least matches or exceeds the performance of existing modality-specific caching methods.
Supplementary Material
Figure 8: L1 Relative Error Curves of different architecture components for DiT-XL. Curves are plotted with 95% confidence intervals and scaled to the same y-axis range.
6 Ablations Addendum: Number of Calibration Samples
As mentioned in Sec. 3.3, the calibration sample size chosen does not dramatically change the generated schedule. Increasing the number of samples only affects the range of the confidence interval for the error curves, but not the mean. Future methods that might leverage the uneven distribution of error curves between samples might need to take note of this observation.